# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ 2.1 Incorporating heterogeneity
    $\begin{pmatrix}\int_{\Omega^{e(1)}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \int_{\Omega^{e(1)}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \int_{\Omega^{e(1)}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\end{pmatrix}\boldsymbol{\theta}=\begin{pmatrix}R^{1}_{i}\\ \vdots\\ R^{k}_{i}\\ \vdots\\ R^{n_{b}}_{i}\end{pmatrix}\quad\forall\quad\begin{split}i&=1,2,3\end{split}$ (18)
where $\Omega_{k}^{e(j)}$ denotes the volume enclosed by elements of material segment $j$ in the volume $\Omega_{k}^{e}$, and $\boldsymbol{\theta}$ retains its definition from Equation 15. Equation 18 results in a system of $3n_{b}$ linear equations following the numerical integration and thus can be assembled as follows
$\boldsymbol{A}^{\text{fix}}\boldsymbol{\theta}=\boldsymbol{b}^{\text{fix}}\quad\text{with}\quad\boldsymbol{A}^{\text{fix}}\in\mathbb{R}^{3n_{b}\times n_{f}n_{c}}\quad\text{and}\quad\boldsymbol{b}^{\text{fix}}\in\mathbb{R}^{3n_{b}}$ (19)
@@
By combining Equation 15 and 19, we formulate a unified model of the form $\boldsymbol{A\theta}=\boldsymbol{b}$, where a hyperparameter $\lambda_{r}$ is introduced to control the relative weights between the free and fixed degrees of freedom in the reference domain within the force balance equation.
@@
Figure 3: Graphical illustration of the unified global heterogenized model $\boldsymbol{A\theta}=\boldsymbol{b}$. The matrix $\boldsymbol{A}$ comprises $\mathcal{D}^{\text{free}}$ rows corresponding to the free degrees of freedom, along with additional rows for the Dirichlet boundary nodes $\mathcal{D}^{\text{fix}}$. Each segment-specific block $\boldsymbol{A}^{(j)}$, containing $n_{f}$ columns associated with the constitutive feature library $\boldsymbol{Q}$, is zero-padded (gray regions) to ensure dimensional compatibility. It is assembled via column-wise concatenation over $n_{c}$ identified material segments. The parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{\mathit{n_{f}n_{c}}\times 1}$ encodes the model coefficients for each material segment. The right-hand side vector $\boldsymbol{b}$ has two components: $\boldsymbol{b}^{\text{free}}=\boldsymbol{0}$ representing equilibrium at internal nodes, and $\boldsymbol{b}^{\text{fix}}$ containing the boundary-aggregated reaction force.
@@
To ensure dimensional compatibility, zero-padding is applied to the rows (d.o.f.) of $\boldsymbol{A}^{(j)}$ not belonging to the material segment $j$. The resulting right-hand side vector $\boldsymbol{b}$ consists of two components: $\boldsymbol{b}^{\text{free}}$, representing the zero internal force condition at nodes in $\mathcal{D}^{\text{free}}$, and $\boldsymbol{b}^{\text{fix}}$, denoting the boundary-aggregated reaction forces $R^{k}_{i}$ $\left(k\in\{1,\dots,n_{b}\};i\in\{1,2,3\}\right)$ for the subset $\mathcal{D}^{\text{fix}}$.
Thus, as shown in Figure 3, the problem of obtaining each segment’s material parameters $\boldsymbol{\theta}^{(j)}$, can be recast into the form $\boldsymbol{A\theta}=\boldsymbol{b}$. From the shape of $\boldsymbol{A}$, it can be noted that $\left|\mathcal{D}^{\text{free}}\right|+3n_{b}>>n_{f}n_{c}$. This would imply $\boldsymbol{A\theta}=\boldsymbol{b}$ is an overdetermined system, and the solution for $\boldsymbol{\theta}$ can be obtained in an ordinary least-square sense as
@@
However, such a solution would suffer from two main drawbacks – (i) the solution would be susceptible to overfitting and displacement noise [38] (see supplementary subsection S1.2), (ii) we would be unable to ensure that each component of $\boldsymbol{\theta}$ would be non-negative, which is essential to ensure polyconvexity of the strain energy density [46]. Thus, following [46] and [54], we formulate the equilibrium problem as a sparse-prior Bayesian regression for $\boldsymbol{\theta}$. The spike-slab prior [54] is used for $\boldsymbol{\theta}$ to enforce parsimony, avoid overfitting, and provide noise-robustness.
@@
Equation 22 forms a likelihood probability, and using Bayes’ rule we can extract the posterior distribution $p(\boldsymbol{\theta},\boldsymbol{z},p_{0},\nu_{s},\sigma^{2}\>|\>\boldsymbol{A},\boldsymbol{b})$ using the approach outlined in Joshi et al. [46]. For brevity, we do not reiterate the Bayesian-EUCLID approach in the main text. A synopsis of the approach is provided in A. The solution for the material models $\boldsymbol{\theta}_{(k)}\left(k\in\{1\dots N_{G}\}\right)$ would be a collection of $N_{G}$ number of values corresponding to the retained elements of the Markov chain, which approximate the posterior distribution for $\boldsymbol{\theta}$. For practical purposes, such as use in forward finite element simulations, it would be convenient to use the approximated mean of the distribution $\boldsymbol{\theta}_{\text{mean}}=\dfrac{\sum_{k\in\{1\dots N_{G}\}}\boldsymbol{\theta}_{(k)}}{N_{G}}$ as material model parameters. The standard deviation $\boldsymbol{\theta}_{\text{std}}=\sqrt{\dfrac{\sum_{k\in\{1\dots N_{G}\}}\left(\boldsymbol{\theta}_{(k)}-\boldsymbol{\theta}_{\text{mean}}\right)^{2}}{N_{G}}}$ provides an approximate measure of the uncertainty in the prediction. The distribution for $\boldsymbol{\theta}$ is multivariate, and $\boldsymbol{\theta}_{\text{std}}$ provides only a practical, but approximate measure of the uncertainty associated with the material models. In this study, we provide the results as $\boldsymbol{\theta}_{\text{mean}}\pm\boldsymbol{\theta}_{\text{std}}$, instead of the violin plots in Joshi et al. [46].
@@
Figure 4: Comparative study of different segmentation techniques. (a) shows the ground truth for a heterogeneous bi-material specimen consisting of a cross-shaped inclusion governed by the Isihara (ISH) constitutive model, embedded in a base matrix described by the Neo-Hookean (NH2) model. (b) presents the segmentation result using $k$-means clustering based on strain invariants [55]. (c) utilizes the complete material feature library (Equation 6) to compute the 2D residual force norm distribution $(f^{a}_{\text{res}})$, and segments the material region by flagging nodes with residual force values above a specified threshold. (d) illustrates the result from our proposed method, where segmentation is performed using a reduced Neo-Hookean feature set $(\tilde{I}_{1}-3)$ and $(J-1)^{2}$, yielding clean and accurate material cluster identification in strong agreement with the ground truth.
@@ 2.2 Segmenting based on displacement data
Nguyen and Lejeune [55] provide a study on segmenting mechanically heterogeneous hyperelastic materials using surface displacement data. The isotropic invariants of the Cauchy-Green tensor ($I_{1},I_{2}$ and $I_{3}$), alongside the displacement vector at the centroid of each finite element, were used as features to cluster each element using either k-means clustering or spectral clustering. This approach was validated on various heterogeneity patterns- circular inclusions, cross-shaped inclusions, Cahn-Hillard patterns [56], and split domains. In most cases, their results found that the segmented regions approximated the underlying true segments and contained non-smooth boundaries. Furthermore, their results did not account for any correlated or uncorrelated noise in the displacement field, which could be expected to deteriorate the segmentation further. Subpart (b) of Figure 4 provides results of k-means clustering of strain invariants and displacements from a non-equi-biaxial tension test on a square geometry with a cross-shaped inclusion. In Figure 4b, it can be seen that a significant number of elements are mis-clustered. B demonstrates that even 25 mis-clustered elements can deteriorate material property prediction accuracy by around 30%. Thus, any adopted segmentation approach should provide accurate boundaries in order to achieve reasonably accurate material property prediction.
@@
In this work, we adopt two steps to achieve segmentation – the first step is to identify the inter-segment boundaries in the unknown material, and the next step is to use the identified boundaries to register all elements with their respective segment. The first step – identification of the inter-segment boundaries, is based on using “residual forces” within the geometry for a homogeneous model assumption. The $\boldsymbol{A}$ (Equation 20) matrix is initially constructed assuming that the geometry contains only one segment (homogeneous) and only the NH2 features- $(\tilde{I}_{1}-3)$ and $(J-1)^{2}$ in the library. This is done only for segmenting the heterogeneous material to get a sharper inter-material interface (see Figure 4d), and in the subsequent Bayesian-EUCLID solution step, we keep all features in the library active. The homogeneous matrix $\boldsymbol{A}$ constructed here is denoted by $\boldsymbol{A}_{0}(\in\mathbb{R}^{(|\mathcal{D}^{\text{free}}|+3n_{b})\times 2})$ to disambiguate it from the fully segmented $\boldsymbol{A}$ matrix that will be used in the Bayesian-EUCLID step. As seen in the discussion preceding Equation 15, $\boldsymbol{b}^{\text{free}}$ indicates the net force acting on each node within the volume of the material. For an appropriate guess of non-trivial material parameters ($\boldsymbol{\theta^{*}}$), we would have $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta}^{*}=\boldsymbol{b}^{\text{free}}=\boldsymbol{0}$. However, no non-trivial solution exists for $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta}=\boldsymbol{0}$. The Ordinary-Least Square (OLS) approximation $\boldsymbol{\theta^{\text{ols}}}$ (Equation 21) leaves the residue $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta^{\text{ols}}}=\boldsymbol{b^{\text{residue}}}\neq\boldsymbol{b^{\text{free}}}$. Thus, we compute $\boldsymbol{b^{\text{residue}}}$ for the homogeneous material system $\boldsymbol{A}_{0}$ as follows
@@
Figure 6: (a)-(f) Schematic illustration of the random seed-based growth algorithm for segmenting material sub-domains. The algorithm initiates from a randomly selected seed element (wedge type) and iteratively grows by adding neighboring elements until it reaches flagged interface boundaries. A new seed element is then chosen from the remaining unvisited elements, and the growth process is repeated until all distinct regions are identified, terminating at either interface or external boundaries.
@@
Nodes with the 2D force magnitude $f^{a}_{\text{res}}$, higher than a certain threshold are “flagged” as inter-segment boundary nodes. Generally, the magnitude of residual force is commensurate with the extent of mismatch between the “guessed” material property (homogeneous assumption) and the true underlying material property. At the nodes that lie on the interface of two (or more) material segments, a homogeneous material assumption would induce significantly more residual force due to the imbalance between elements with different material properties. Thus, these interface nodes would be characterized by a significantly higher than average (across all nodes) residual force. Therefore, we employ the following scheme to “flag” inter-segment nodes.
$a\in\boldsymbol{{N}_{\text{flag}}},\quad\iff\quad f_{\text{res}}^{a}>\lambda\,\sigma_{f_{\text{res}}}$ (26)
where $a$ is an integer, $\boldsymbol{{N}_{\text{flag}}}$ denotes the set of all flagged nodes, $\lambda$ is a tunable scalar (usually ranging between 1.5 and 2.5) and the standard deviation $\sigma_{f_{\text{res}}}$ is given by
(a) Type 1: Cross-shaped inclusion
(b) Type 2: Split-domain pattern
(c) Type 3: Multiple inner inclusions
(d) Type 4: Cahn-Hilliard pattern
Figure 7: Graphical illustration of heterogeneous test specimen geometries subjected to non-equi-biaxial tensile loading under three-dimensional plane stress conditions. The heterogeneity patterns include: (a) cross-shaped inclusion, (b) split-domain configuration, (c) multiple non-circular inclusions, and (d) Cahn-Hilliard pattern.
$\sigma_{f_{\text{res}}}=\sqrt{\frac{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}\left(f^{a}_{\text{res}}-\mu\right)^{2}}{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}1}}\quad;\quad\mu_{f_{\text{res}}}=\frac{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}\left(f^{a}_{\text{res}}\right)}{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}1}$
For zero and moderate noise, it was observed that the average of $f^{a}_{\text{res}}$ ,i.e, $\mu_{f_{\text{res}}}$ was significantly smaller than its standard deviation $\sigma_{f_{\text{res}}}$. However, $\mu_{f_{\text{res}}}$ was observed to increase with an increase in displacement noise, while $\sigma_{f_{\text{res}}}$ was nearly constant. This indicates that $\mu_{f_{\text{res}}}$ is a measure of the noise prevalent in the displacement measurement. Further, this implies that $f^{a}_{\text{res}}$ could be high in nodes that are either on the inter-segment boundary or have high displacement noise. In order to disambiguate between the inter-segment nodes and those with high displacement noise, we employ a scalar multiple of $\sigma_{f_{\text{res}}}$ as a lower threshold. In experiments where the displacement noise is sufficiently low (measured by the ratio: $\mu_{f_{\text{res}}}/\sigma_{f_{\text{res}}}$), the inter-segment nodes can be expected to mostly have $f^{a}_{\text{res}}$ above the lower threshold, while the “noisy” nodes would mostly have $f^{a}_{\text{res}}$ below the threshold. This distinction blurs with increasing displacement noise, making residual force-based segmentation inaccurate for data with high displacement noise. In subsection 3.4 we provide further discussion on the observed statistical trends for nodal residual force $f^{a}_{\text{res}}$. Notably, Shi et al. [45] makes similar use of residual forces from a homogenized assumption of their NN surrogate model to highlight the inter-subdomain boundaries, with user-defined lower thresholds.
@@
2:Definitions: $N_{a}\rightarrow$ Node $a$, with $a\in\{1\dots n_{\text{nodes}}\}$
@@
33:i1i-1 indicates the number of discovered material segments

### v2

@@ 2.1 Incorporating heterogeneity
    $\begin{pmatrix}\int_{\Omega^{e(1)}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \int_{\Omega^{e(1)}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \int_{\Omega^{e(1)}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\int_{\Omega^{e(n_{c})}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\end{pmatrix}\boldsymbol{\theta}=\begin{pmatrix}R^{1}_{i}\\ \vdots\\ R^{k}_{i}\\ \vdots\\ R^{n_{b}}_{i}\end{pmatrix}\quad\text{where}\quad\begin{split}i=1,2,3\end{split}$ (18)
where $\Omega_{k}^{e(j)}$ denotes the volume enclosed by elements of material segment $j$ in the volume $\Omega_{k}^{e}$, and $\boldsymbol{\theta}$ retains its definition from Equation 15. However, in routine material testing experiments that do not use specialized shear beam load cells, the load cells measure only one component of force, i.e, normal force. Designating the $i^{\text{th}}$ component of the normal vector and magnitude of the normal force at the $k^{\text{th}}$ boundary as $n^{k}_{i}$ and $F^{k}$, respectively, we have,
$F^{k}=\sum_{i=1}^{3}n^{k}_{i}R_{i}^{k}\qquad k=1,2,\dots,n_{b}$ (19)
Thus, Equation 18 gets modified to:
$\begin{pmatrix}\sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(1)}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(n_{c})}_{1}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(1)}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(n_{c})}_{k}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\\ \vdots\\ \sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(1)}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\quad\dots\quad\sum_{i=1}^{3}n_{i}^{k}\int_{\Omega^{e(n_{c})}_{n_{b}}}\frac{\partial\boldsymbol{Q}^{T}}{\partial F_{i\mathrm{J}}}\mathbb{N}_{,\mathrm{J}}^{a}(\boldsymbol{X})\mathrm{~d}V\end{pmatrix}\boldsymbol{\theta}=\begin{pmatrix}F^{1}\\ \vdots\\ F^{k}\\ \vdots\\ F^{n_{b}}\end{pmatrix}$ (20)
Equation 20 results in a system of $n_{b}$ linear equations (one corresponding to each load-cell reading) following numerical integration and thus can be assembled as follows
$\boldsymbol{A}^{\text{fix}}\boldsymbol{\theta}=\boldsymbol{b}^{\text{fix}}\quad\text{with}\quad\boldsymbol{A}^{\text{fix}}\in\mathbb{R}^{n_{b}\times n_{f}n_{c}}\quad\text{and}\quad\boldsymbol{b}^{\text{fix}}\in\mathbb{R}^{n_{b}}$ (21)
@@
By combining Equation 15 and 21, we formulate a unified model of the form $\boldsymbol{A\theta}=\boldsymbol{b}$, where a hyperparameter $\lambda_{r}$ is introduced to control the relative weights between the free and fixed degrees of freedom in the reference domain within the force balance equation.
@@
Figure 3: Graphical illustration of the unified global heterogenized model $\boldsymbol{A\theta}=\boldsymbol{b}$. The matrix $\boldsymbol{A}$ comprises $\mathcal{D}^{\text{free}}$ rows corresponding to the free degrees of freedom, along with additional rows for the Dirichlet boundary nodes $\mathcal{D}^{\text{fix}}$. Each segment-specific block $\boldsymbol{A}^{(j)}$, containing $\mathit{n_{f}}$ columns associated with the constitutive feature library $\boldsymbol{Q}$, is zero-padded (gray regions) to ensure dimensional compatibility. It is assembled via column-wise concatenation over $n_{c}$ identified material segments. The parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{\mathit{n_{f}n_{c}}\times 1}$ encodes the model coefficients for each material segment. The right-hand side vector $\boldsymbol{b}$ has two components: $\boldsymbol{b}^{\text{free}}=\boldsymbol{0}$ representing equilibrium at internal nodes, and $\boldsymbol{b}^{\text{fix}}$ containing the boundary-aggregated reaction force.
@@
To ensure dimensional compatibility, zero-padding is applied to the rows (d.o.f.) of $\boldsymbol{A}^{(j)}$ not belonging to the material segment $j$. The resulting right-hand side vector $\boldsymbol{b}$ consists of two components: $\boldsymbol{b}^{\text{free}}$, representing the zero internal force condition at nodes in $\mathcal{D}^{\text{free}}$, and $\boldsymbol{b}^{\text{fix}}$, denoting the boundary-aggregated reaction forces $R^{k}_{i}$ $\left(k\in\{1,\dots,n_{b}\}\,;\,i(k)=1\text{ or }2\right)$ for the subset $\mathcal{D}^{\text{fix}}$.
Thus, as shown in Figure 3, the problem of obtaining each segment’s material parameters $\boldsymbol{\theta}^{(j)}$, can be recast into the form $\boldsymbol{A\theta}=\boldsymbol{b}$. From the shape of $\boldsymbol{A}$, it can be noted that $\left|\mathcal{D}^{\text{free}}\right|+n_{b}>>n_{f}n_{c}$. This would imply $\boldsymbol{A\theta}=\boldsymbol{b}$ is an overdetermined system, and the solution for $\boldsymbol{\theta}$ can be obtained in an ordinary least-square sense as
@@
However, such a solution would suffer from two main drawbacks – (i) the solution would be susceptible to overfitting and displacement noise [Flaschel et al., 2021a] (see supplementary subsection S1.2), (ii) we would be unable to ensure that each component of $\boldsymbol{\theta}$ would be non-negative, which is essential to ensure polyconvexity of the strain energy density [Joshi et al., 2022]. Thus, following [Joshi et al., 2022] and [Nayek et al., 2021], we formulate the equilibrium problem as a sparse-prior Bayesian regression for $\boldsymbol{\theta}$. The spike-slab prior [Nayek et al., 2021] is used for $\boldsymbol{\theta}$ to enforce parsimony, avoid overfitting, and provide noise-robustness.
@@
Equation 24 forms a likelihood probability, and using Bayes’ rule, we can extract the posterior distribution $p(\boldsymbol{\theta},\boldsymbol{z},p_{0},\nu_{s},\sigma^{2}\>|\>\boldsymbol{A},\boldsymbol{b})$ using the approach outlined in Joshi et al. [2022]. For brevity, we do not reiterate the Bayesian-EUCLID approach in the main text. A synopsis of the approach is provided in A. The solution for the material models $\boldsymbol{\theta}_{(k)}\left(k\in\{1\dots N_{G}\}\right)$ would be a collection of $N_{G}$ number of values corresponding to the retained elements of the Markov chain, which approximate the posterior distribution for $\boldsymbol{\theta}$. For practical purposes, such as use in forward finite element simulations, it would be convenient to use the approximated mean of the distribution $\boldsymbol{\theta}_{\text{mean}}=\dfrac{\sum_{k\in\{1\dots N_{G}\}}\boldsymbol{\theta}_{(k)}}{N_{G}}$ as material model parameters. The standard deviation $\boldsymbol{\theta}_{\text{std}}=\sqrt{\dfrac{\sum_{k\in\{1\dots N_{G}\}}\left(\boldsymbol{\theta}_{(k)}-\boldsymbol{\theta}_{\text{mean}}\right)^{2}}{N_{G}}}$ provides an approximate measure of the uncertainty in the prediction. The distribution for $\boldsymbol{\theta}$ is multivariate, and $\boldsymbol{\theta}_{\text{std}}$ provides only a practical, but approximate measure of the uncertainty associated with the material models. In this study, we provide the results as $\boldsymbol{\theta}_{\text{mean}}\pm\boldsymbol{\theta}_{\text{std}}$, instead of the violin plots in Joshi et al. [2022].
2.2 Comment on Equation 24
An important assumption that is made in writing Equation 24 is that the “noise” term $\boldsymbol{\epsilon}$ can be sampled from a normal distribution with a diagonal covariance matrix. Since Equation 24 forms a likelihood of satisfying the weak form of the linear momentum balance, the term $\boldsymbol{A\theta-b}$ can be interpreted as a nodal residual force $\boldsymbol{f}_{\mathrm{res}}$. Thus, $\boldsymbol{\epsilon}$ serves as a “noise” for the nodal residual force. However, as shown by Avril and Pierron [2007] for the deformation of a linear elastic solid, a spatially uncorrelated displacement noise induces a spatially correlated noise in the nodal residual force $\boldsymbol{f}_{\mathrm{res}}$($\boldsymbol{A\theta-b}$). This implies that $\boldsymbol{\epsilon}$ cannot be sampled from a normal distribution with a diagonal covariance matrix. In B we perform spatial correlation studies on the residual force vector field $\boldsymbol{f}_{\mathrm{res}}$ for a homogeneous hyperelastic material to verify that for a given displacement field, the nodal residual forces are indeed spatially correlated, i.e, possess a non-diagonal covariance matrix. Further, in C we demonstrate that a similar incorrect assumption regarding the lack of spatial correlation in nodal residual forces is made by many important contemporary stress-unsupervised constitutive modeling frameworks that use the mean-squared error loss function [Thakolkaran et al., 2022, Shi et al., 2025, Meng et al., 2025] or use a variation of the ordinary least square solution to determine the material parameters [Haustrate et al., 2024, Flaschel et al., 2021a]. However, as will be seen in section 3, the mean value of the material parameters ($\boldsymbol{\theta}_{\text{mean}}$) obtained from sampling the posterior probability distribution are predicted with good accuracy, despite the incorrect assumption made in Equation 24. By accounting for spatial correlations in the residual force vector in Equation 24, we anticipate more robust uncertainty quantification of the predicted material parameters. In the current work, we proceed with the use of Equation 24 as the likelihood distribution to obtain the posterior material parameter distribution, without accounting for spatial correlations in the nodal residual force.
@@ 2.2 Comment on Equation 24
Figure 4: Comparative study of different segmentation techniques. (a) shows the ground truth for a heterogeneous bi-material specimen consisting of a cross-shaped inclusion governed by the Isihara (ISH) constitutive model, embedded in a base matrix described by the Neo-Hookean (NH2) model. (b) presents the segmentation result using $k$-means clustering based on strain invariants [Nguyen and Lejeune, 2024b]. (c) utilizes the complete material feature library (Equation 6) to compute the 2D residual force norm distribution $(f^{a}_{\text{res}})$, and segments the material region by flagging nodes with residual force values above a specified threshold. (d) illustrates the result from our proposed method, where segmentation is performed using a reduced Neo-Hookean feature set $(\tilde{I}_{1}-3)$ and $(J-1)^{2}$, yielding clean and accurate material cluster identification in strong agreement with the ground truth.
@@ 2.3 Segmenting based on displacement data
Nguyen and Lejeune [2024b] provide a study on segmenting mechanically heterogeneous hyperelastic materials using surface displacement data. The isotropic invariants of the Cauchy-Green tensor ($I_{1},I_{2}$ and $I_{3}$), alongside the displacement vector at the centroid of each finite element, were used as features to cluster each element using either k-means clustering or spectral clustering. This approach was validated against various heterogeneity patterns, including circular inclusions, cross-shaped inclusions, Cahn-Hillard patterns [Kobeissi et al., 2022], and split domains. In most cases, their results found that the segmented regions approximated the underlying true segments and contained non-smooth boundaries. Furthermore, their results did not account for correlated or uncorrelated noise in the displacement field, which could be expected to deteriorate the segmentation further. Subpart (b) of Figure 4 provides results of k-means clustering of strain invariants and displacements from a non-equi-biaxial tension test on a square geometry with a cross-shaped inclusion. In Figure 4b, it can be seen that a significant number of elements are mis-clustered. D demonstrates that even 25 mis-clustered elements can deteriorate material property prediction accuracy by around 30%. Thus, any adopted segmentation approach should provide accurate boundaries in order to achieve reasonably accurate material property predictions.
@@
In this work, we adopt a two-step approach to achieve segmentation. The first step involves identifying the inter-segment boundaries in the unknown material, and the second step utilizes these identified boundaries to register all elements with their respective segments. The identification of inter-segment boundaries is based on using “residual forces” within the geometry, assuming a homogeneous model. The $\boldsymbol{A}$ (Equation 22) matrix is initially constructed assuming that the geometry contains only one segment (homogeneous) and only the NH2 features, $(\tilde{I}_{1}-3)$ and $(J-1)^{2}$ in the library. This is done only for segmenting the heterogeneous material to get a sharper inter-material interface (see Figure 4d), and in the subsequent Bayesian-EUCLID solution step, we keep all features in the library active. The homogeneous matrix $\boldsymbol{A}$ constructed here is denoted by $\boldsymbol{A}_{0}(\in\mathbb{R}^{(|\mathcal{D}^{\text{free}}|+n_{b})\times 2})$ to disambiguate it from the fully segmented $\boldsymbol{A}$ matrix that will be used in the Bayesian-EUCLID step. As seen in the discussion preceding Equation 15, $\boldsymbol{b}^{\text{free}}$ indicates the net force acting on each node within the volume of the material. For an appropriate guess of non-trivial material parameters ($\boldsymbol{\theta^{*}}$), we would have $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta}^{*}=\boldsymbol{b}^{\text{free}}=\boldsymbol{0}$. However, no non-trivial solution exists for $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta}=\boldsymbol{0}$. The Ordinary-Least Square (OLS) approximation $\boldsymbol{\theta^{\text{ols}}}$ (Equation 23) leaves the residue $\boldsymbol{A^{\text{free}}}\boldsymbol{\theta^{\text{ols}}}=\boldsymbol{b^{\text{residue}}}\neq\boldsymbol{b^{\text{free}}}$. Thus, we compute $\boldsymbol{b^{\text{residue}}}$ for the homogeneous material system $\boldsymbol{A}_{0}$ as follows
@@
Figure 6: (a)-(f) Schematic illustration of the random seed-based growth algorithm for segmenting material sub-domains. The algorithm initiates from a randomly selected seed element (wedge type) and iteratively grows by adding neighboring elements until it reaches flagged interface boundaries. A new seed element is then chosen from the remaining unvisited elements, and the growth process is repeated until all distinct regions are identified, terminating at either an interface or an external boundary.
@@
The nodes with the 2D force magnitude $f^{a}_{\text{res}}$, higher than a certain threshold, are “flagged” as inter-segment boundary nodes. Generally, the magnitude of residual force is commensurate with the extent of mismatch between the “guessed” material property (homogeneous assumption) and the underlying true material property. At the nodes that lie on the interface of two (or more) material segments, a homogeneous material assumption would induce significantly larger residual force due to the imbalance between elements with different material properties. Thus, these interface nodes would be characterized by a considerably higher than average (across all nodes) residual force. Therefore, we employ the following scheme to “flag” inter-segment nodes.
$a\in\boldsymbol{{N}_{\text{flag}}},\quad\iff\quad f_{\text{res}}^{a}>\lambda\,\mathit{\sigma_{f_{\text{res}}}}$ (28)
where $a$ is an integer, $\boldsymbol{{N}_{\text{flag}}}$ denotes the set of all flagged nodes, $\lambda$ is a tunable scalar (usually ranging between 1.5 and 2.5) and the standard deviation $\mathit{\sigma_{f_{\text{res}}}}$ is given by
@@
2:Definitions: $N_{a}\rightarrow$ Node $a$, with $a\in\{1\dots n_{n}\}$
@@
33:  Initialize $\mathcal{S}_{0}\leftarrow\mathcal{S}_{\mathrm{domain}}\setminus\left(\cup_{k=1}^{i}\mathcal{S}_{k}\right)$